Josh can row a boat at an average speed that is eight times as fast as his average swimming speed. It takes him 80 minutes more to swim 4 laps in a lake than to row 4 laps. How fast is his average swimming speed in laps per minute? Round your answer to two decimal places.
Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)

Josh can row a boat at an average speed that is eight times as fast as his average swimming speed. It takes him 80 minutes more to swim 4 laps in a lake than to row 4 laps. How fast is his average swimming speed in laps per minute? Round your answer to two decimal places.
------
Swim DATA:
rate = r laps/min ; time = (t+80) min ; distance = r*t + 80r laps
---------------
Row DATA:
rate = 8r laps/min ; time = t min ; distance = 8rt laps
-------------------------
Equation:
rt+80r = 8rt
7rt = 80r
time = 80/7 = 11.42 min for 4 laps
----
Ans: 11.42/4 = 2.855 min per lap
rate = 1/2.855 = 0.35 laps per min
---------------
Cheers,
Stan H.
----------------


Answer by MathTherapy(10555)


Josh can row a boat at an average speed that is eight times as fast as his average swimming speed. It takes him 80 minutes more to swim 4 laps in a lake than to row 4 laps. How fast is his average swimming speed in laps per minute? Round your answer to two decimal places.